# Stylometry recognizes human and LLM-generated texts in short samples

### v1

@@ Abstract
The paper explores stylometry as a method to distinguish between texts created by Large Language Models (LLMs) and humans, addressing issues of model attribution, intellectual property, and ethical AI use. Stylometry has been used extensively to characterise the style and attribute authorship of texts. By applying it to LLM-generated texts, we identify their emergent writing patterns. The paper involves creating a dataset based on Wikipedia, processed through multiple text summarization methods (T5, BART, Gensim, and Sumy) and LLMs (GPT-3.5, GPT-4, LLaMa 2/3, Orca, and Falcon). The 10-sentence long texts were classified by tree-based models (decision trees and LightGBM) using human-designed (StyloMetrix) and n-gram-based (our own pipeline) stylometric features that encode lexical, grammatical, syntactic, and punctuation patterns. The cross-validated results reached a performance of up to .87 Matthews correlation coefficient in the multiclass scenario with 7 classes, and accuracy between .79 and 1. in binary classification, with the particular example of Wikipedia vs GPT-4 reaching up to .98 accuracy on a balanced data set. Shapley Additive Explanations pinpoint features characteristic of the encyclopaedic text type, individual overused words, as well as a greater grammatical standardisation of LLMs with respect to human-written texts. These results show – crucially, in the context of the increasingly sophisticated LLMs, like GPT-3.5 and GPT-4 – that it is possible to distinguish machine- from human-generated texts at least for a well-defined text type. We emphasise the need for robust techniques to track AI outputs and ensure ethical use.
@@
stylometry, large language models
@@ 1 Introduction
In the rapidly developing landscape of natural language processing, Large Language Models (LLMs), delineated by transformative models like GPT-3.5, have revolutionized natural language processing, enabling machines to mimic human-like text generation. As pretrained models become more prevalent, concerns regarding model ownership and attribution, intellectual property, and responsible AI utilization underscore the importance of developing advanced techniques to ensure ethical use and proper attribution of Artificial Intelligence-generated content and the need for effective model detection techniques.
The problem of stylometry and authorship attribution emerges as a crucial aspect in this context. Stylometry, the quantitative study of linguistic style patterns, serves as a valuable asset for effective text differentiation. By examining subtle variations in writing style, one can unveil unique markers that distinguish one author from another. Stylometric features provide a nuanced understanding of the individual characteristics of LLMs, offering a granular approach to model identification. This not only facilitates differentiation but also enhances our comprehension of the linguistic idiosyncrasies ingrained in these models. The challenge lies in accurately attributing text to the correct author or model, especially as language models grow more sophisticated and their outputs increasingly indistinguishable from human writing. In the present paper, we train machine learning models to detect subtle stylistic features and patterns characteristic of specific language models, enabling their more precise differentiation. By analyzing distinctive writing style features, such as vocabulary choices and sentence structures, we aim to shed light on the linguistic nuances that set each model apart.
The exploration of stylometry in model detection and differentiation extends beyond technical considerations to ethical implications. Understanding the distinct stylometric features of language models contributes to responsible AI practices, promoting transparency and accountability in their deployment. LLM safety and ethics are paramount concerns in this regard. Ensuring that language models are used ethically involves addressing issues such as bias, misinformation, and the potential for generating harmful content. By embracing stylometry, this paper aims to provide a nuanced perspective , thereby contributing to a more comprehensive understanding of language model deployment in diverse applications. This approach not only enhances our ability to safeguard intellectual property but also fosters a culture of responsibility and trust in the AI community.
@@
    Application of Stylometry to Differentiate Texts: The paper applies stylometry to distinguish between texts generated by Large Language Models (LLMs) and human-authored texts. Stylometry, traditionally used for authorship attribution and literary style analysis, is shown to be effective in identifying writing patterns specific to LLMs.
@@
    Creation of a Diverse Dataset: The study constructs a dataset based on Wikipedia texts and their summaries, processed through various text summarization methods (T5, BART, Gensim, and Sumy) and LLMs (GPT-3.5, GPT-4, LLaMa 2/3, Orca, and Falcon). This dataset allows for a comprehensive analysis of different text generation methods.
@@ 2 Related works
Stylometry, the study of linguistic style, has long been a important tool in authorship attribution, and its relevance has grown significantly with the advent of Large Language Models. As these models produce increasingly human-like text, the ability to distinguish between human-authored and LLM-generated content becomes essential, not just for academic and forensic purposes, but also for ensuring the safety and ethical use of LLMs. The application of stylometry to LLMs is particularly important given the potential risks associated with their misuse, such as the generation of misleading information, deepfake text, or malicious content. In this section we present works relevant to the theme of stylometry itself and related to the LLMs; we mention research about stylometric modeling; and finally showcase papers that tackle the theme of safety and ethics regarding emerging generative linguistic tools.
@@ 2.1 Stylometry and author attribution
(Neal et al., 2017) in Surveying Stylometry Techniques and Applications provide an extensive overview of stylometry research, focusing on authorship attribution, verification, profiling, stylochronometry, and adversarial stylometry. The survey is thorough, covering various subtasks, datasets, experimental methods, and contemporary approaches. It includes a detailed performance analysis involving 1,000 authors using 14 different algorithms. The paper highlights key challenges such as scaling authorship analysis techniques to handle a large number of authors with minimal text samples. It also addresses ongoing research challenges and introduces various software tools that support stylometry tasks, showcasing both open-source and commercial options.
@@ 2.2 Stylometric modeling
Paper titled TDRLM: Stylometric learning for authorship verification by Topic-Debiasing (Hu et al., 2023) proposes a "Topic-Debiasing Representation Learning Model" (TDRLM) to enhance stylometric authorship verification. The TDRLM utilizes a topic-debiasing attention mechanism with position-specific topic scores to mitigate the influence of topical bias in tokenized texts. Experimental results demonstrate that the TDRLM outperforms current state-of-the-art stylometric learning models and advanced language models, achieving the highest Area Under Curve (AUC) scores of 92.47% for the Twitter-Foursquare dataset and 93.11% for the ICWSM Twitter dataset. The study highlights that topic-related words can negatively impact machine learning algorithms for authorship verification, prompting the development of the TDRLM model to improve verification accuracy.
In the Neural Authorship Attribution: Stylometric Analysis on Large Language Models Kumarage & Liu (2023) explore methods to identify the source of AI-generated text by examining unique writing signatures of different LLMs. The study distinguishes between proprietary models (like GPT-4 and GPT-3.5) and open-source models (such as Llama and GPT-NeoX), using a combination of stylometric features – lexical, syntactic, and structural – to enhance neural authorship attribution. The findings indicate that while distinct writing styles can differentiate LLMs, advancements in open-source models may narrow these distinctions, posing challenges for future authorship attribution efforts.
@@ 2.3 Authorship-stylometry and LLMs
Learning Stylometric Representations for Authorship Analysis (Ding et al., 2017) published in IEEE Transactions on Cybernetics, explores a neural network approach to learn stylometric representations that capture various linguistic features such as topical, lexical, syntactical, and character-level characteristics. This methodology aims to improve the tasks of authorship characterization, identification, and verification by mimicking the human sentence composition process and incorporating these diverse linguistic categories into a distributed representation of words. The effectiveness of this approach is demonstrated through extensive evaluations across multiple datasets, including Twitter, blogs, reviews, novels, and essays, where the proposed models notably outperform traditional stylometric and other baseline methods. This research highlights the potential of neural networks in extracting and utilizing complex stylistic features for detailed authorship analysis in diverse textual domains.
@@
A model-independent redundancy measure for human versus ChatGPT authorship discrimination using a Bayesian probabilistic approach (Bozza et al., 2023) introduces a novel method to distinguish between human-authored texts and those generated by AI models like ChatGPT. This approach utilizes a model-independent redundancy measure that effectively captures syntactical differences between human and AI-generated texts. The researchers employed a Bayesian probabilistic framework, specifically using the Bayes factor, to provide a robust and consistent classification criterion. This method proves particularly effective even with short text samples, demonstrating its potential utility in forensic and other analytical settings where distinguishing between human and AI authorship is crucial. The study highlights the applicability of this technique across various languages and text genres, indicating its broad potential for addressing the challenges posed by the increasing sophistication of AI-generated text in academic and professional contexts.
@@
Authors of Who Wrote it and Why? Prompting Large Language Models for Authorship Verification (Hung et al., 2023) offer a new technique named PromptAV. This method utilizes Large Language Models (LLMs) to perform authorship verification effectively and with improved interpretability. Authors claim that the PromptAV, demonstrates improved performance compared to existing state-of-the-art baselines, particularly in scenarios with limited training data. It enhances interpretability by providing intuitive explanations, making it a promising tool for applications in forensic analysis, plagiarism detection, and identifying deceptive content in texts. This approach is meant to address the current limitations of traditional stylometric and deep learning methods, which typically require extensive data and lack explainability.
@@
The paper T5 meets Tybalt: Author Attribution in Early Modern English Drama Using Large Language Models (Hicke & Mimno, 2023) explores the application of LLMs for authorship identification in Early Modern English drama. The study finds that LLMs, specifically a fine-tuned T5-large model, can accurately predict the author of short passages and outperform traditional baselines like logistic regression, SVM with a linear kernel, and cosine delta. However, the presence of certain authors in the model’s pre-training data introduces biases, leading to occasional confident misattributions of texts. This highlights both the promising potential and the concerning limitations of using LLMs for stylometric analysis in literary studies.
@@
Finally, the paper titled Detecting ChatGPT: A Survey of the State of Detecting ChatGPT-Generated Text (Dhaini et al., 2023) provides an overview of current approaches for identifying text generated by ChatGPT. It highlights the challenges of distinguishing between human-written and AI-generated content, especially given the high fluency and human-like quality of ChatGPT outputs. The survey reviews various datasets specifically created for this detection task, examines different methodologies employed, and discusses qualitative analyses that help identify characteristics unique to ChatGPT-generated text. It also explores the broader implications for domains such as education, law, and science, emphasizing the need for effective detection methods to maintain content integrity.

### v2

@@ Abstract
The paper explores stylometry as a method to distinguish between texts created by Large Language Models (LLMs) and humans, addressing issues of model attribution, intellectual property, and ethical AI use. Stylometry has been used extensively to characterise the style and attribute authorship of texts. By applying it to LLM-generated texts, we identify their emergent writing patterns. The paper involves creating a benchmark dataset based on Wikipedia, with (a) human-written term summaries, (b) texts generated purely by LLMs (GPT-3.5/4, LLaMa 2/3, Orca, and Falcon), (c) processed through multiple text summarisation methods (T5, BART, Gensim, and Sumy), and (d) rephrasing methods (Dipper, T5). The 10-sentence long texts were classified by tree-based models (decision trees and LightGBM) using human-designed (StyloMetrix) and n-gram-based (our own pipeline) stylometric features that encode lexical, grammatical, syntactic, and punctuation patterns. The cross-validated results reached a performance of up to .87 Matthews correlation coefficient in the multiclass scenario with 7 classes, and accuracy between .79 and 1. in binary classification, with the particular example of Wikipedia and GPT-4 reaching up to .98 accuracy on a balanced dataset. Shapley Additive Explanations pinpointed features characteristic of the encyclopaedic text type, individual overused words, as well as a greater grammatical standardisation of LLMs with respect to human-written texts. These results show – crucially, in the context of the increasingly sophisticated LLMs – that it is possible to distinguish machine- from human-generated texts at least for a well-defined text type.
@@
stylometry, large language models, machine-generated text detection, AI detection, benchmark dataset
@@ 1 Introduction
In the dynamicaly expanding field of natural language processing, Large Language Models (LLMs), introduced by transformative models like GPT, have revolutionized approaches to analyze language by enabling machines to mimic human-like text generation. As the use of pretrained AI models becomes increasingly common, growing concerns around issues like ownership, attribution, intellectual property rights, and responsible usage highlight the urgent need for advanced methods to ensure ethical deployment and proper crediting of AI-generated work – alongside the development of reliable model detection tools.
The problem of stylometry and authorship attribution is a crucial aspect in this context. Stylometry, meaning the quantitative study of linguistic style patterns, is a valuable tool for effective text differentiation. By examining subtle differences in writing style, one can discover unique markers that distinguish one author from another. Stylometric features provide a detailed understanding of the characteristics of particular LLMs, offering a granular approach towards model identification. This not only facilitates differentiation but also enhances our comprehension of the linguistic idiosyncrasies ingrained in these models. The challenge lies in accurately attributing text to the correct author or model, especially as language models grow more sophisticated and their outputs increasingly indistinguishable from human writing. This paper explores the utilization of machine learning techniques in identifying stylistic markers and patterns that are characteristic for specific language models, augmenting our ability to differentiate them with greater accuracy. By focusing on properties such as the word choice and syntactic patterns, our aim is to uncover the linguistic fingerprints that distinguish one model results from another.
The exploration of stylometry in model detection and differentiation reaches far beyond technical considerations towards ethical implications. Understanding the characteristic stylometric properties of language models productions contributes to securing the responsible AI practices, promoting transparency and accountability. LLM safety and ethics are the most important concerns in this regard. Ensuring that language models are used ethically should account for such issues as bias, misinformation, and the potential for generating harmful content. By promoting stylometry, this paper aims to provide a distinctive perspective,thereby contributing to a more comprehensive understanding of language model deployment in diverse applications. This approach not only improves our ability to safeguard intellectual property but also cultivates a culture of responsibility and trust in the AI community.
@@
    Application of Stylometry to Differentiate Texts: The paper applies stylometry to distinguish between texts generated by LLMs and human-authored texts. Stylometry, traditionally used for authorship attribution and literary style analysis, is shown to be effective in identifying writing patterns specific to LLMs.
@@
    Creation of a Diverse Dataset: The study constructs a dataset based on (a) human-written Wikipedia texts, (b) their summaries processed through various text summarization methods (T5, BART, Gensim, and Sumy), and (c) summaries generated by LLMs (GPT-3.5, GPT-4, LLaMa 2/3, Orca, and Falcon) prompted a given term only. This dataset allows for a comprehensive analysis of different text generation methods.
@@ 2 Related works
Stylometry, the study of linguistic style, has long been a important tool in authorship attribution, and its relevance has grown significantly with the advent of Large Language Models. As these models produce increasingly human-like text, the ability to distinguish between human-authored and machine-generated texts (MGT) becomes essential, not just for academic and forensic purposes, but also for ensuring the safety and ethical use of LLMs.
In this section we present works relevant to the theme of stylometry itself and related to MGT, particularly by LLMs; we mention research about stylometric modeling; and finally showcase papers that tackle the theme of safety and ethics regarding emerging generative linguistic tools.
@@ 2.1 Stylometry and author attribution
Neal et al. (2017) in Surveying Stylometry Techniques and Applications provide an extensive overview of stylometry research, focusing on authorship attribution, verification, profiling, stylochronometry, and adversarial stylometry. The survey is in depth, covering different subtasks, datasets, experimental methods, and contemporary approaches. It includes detailed performance analysis taking into account 1,000 authors using 14 different algorithms. The paper exposes key challenges such as scaling authorship analysis to account for a large number of authors with minimal text samples available. It also presents ongoing research challenges and showcases different software tools that support stylometric analysis - both open-source and commercial options.
@@ 2.2 Stylometric modeling
Paper titled TDRLM: Stylometric learning for authorship verification by Topic-Debiasing (Hu et al., 2023) proposes a “Topic-Debiasing Representation Learning Model” (TDRLM) to enhance stylometric authorship verification. The TDRLM utilizes a topic-debiasing attention mechanism with position-specific topic scores to mitigate the influence of topical bias in tokenized texts. Experimental results demonstrate that the TDRLM outperforms current state-of-the-art stylometric learning models and advanced language models, achieving the highest Area Under Curve (AUC) scores of 92.47% for the Twitter-Foursquare dataset and 93.11% for the ICWSM Twitter dataset. The study highlights that topic-related words can negatively impact machine learning algorithms for authorship verification, prompting the development of the TDRLM model to improve verification accuracy.
The evolution of current methods is well exemplified by a series of papers by Kumarage et al. (2023), Kumarage & Liu (2023), Bhattacharjee et al. (2023). Kumarage et al. (2023) and Kumarage & Liu (2023) used a fusion architecture of fine-tuned RoBERTa augmented with a combination of stylometric features – lexical, syntactic, and structural such as lexical richness, readability, punctuation counts, word / sentence / paragraph counts etc. Interestingly, the authors used as a baseline XGBoost with either stylometric or bag-of-word features, which allowed them to use SHAP explanation in the same vein as we do in the present paper. The fusion was proved beneficial especially for short texts (Twitter timelines) and limited training data, but out-of-distribution problems (cross-domain or unseen LLMs) remained challenging. To improve these issues, Bhattacharjee et al. (2023) turned away from stylometric classfiers to self-supervised contrastive learning and unsupervised domain adaptation techniques at the cost of losing the explainability.
@@ 2.3 Authorship-stylometry and LLMs
Learning Stylometric Representations for Authorship Analysis (Ding et al., 2017) explores a neural network approach to learn stylometric representations that capture various linguistic features such as topical, lexical, syntactical, and character-level characteristics. This methodology aims to improve the tasks of authorship characterization, identification, and verification by mimicking the human sentence composition process and incorporating these diverse linguistic categories into a distributed representation of words. The effectiveness of this approach is demonstrated through extensive evaluations across multiple datasets, including Twitter, blogs, reviews, novels, and essays, where the proposed models notably outperform traditional stylometric and other baseline methods. This research highlights the potential of neural networks in extracting and utilizing complex stylistic features for detailed authorship analysis in diverse textual domains.
@@
A model-independent redundancy measure for human versus ChatGPT authorship discrimination using a Bayesian probabilistic approach (Bozza et al., 2023) introduces a novel method to distinguish between human-authored texts and those generated by AI models like ChatGPT. This approach utilizes a model-independent redundancy measure that effectively captures syntactical differences between human and machine-generated texts. The researchers employed a Bayesian probabilistic framework, specifically using the Bayes factor, to provide a robust and consistent classification criterion. This method proves particularly effective even with short text samples, demonstrating its potential utility in forensic and other analytical settings where distinguishing between human and AI authorship is crucial. The study highlights the applicability of this technique across various languages and text genres, indicating its broad potential for addressing the challenges posed by the increasing sophistication of MGT in academic and professional contexts.
@@
Authors of Who Wrote it and Why? Prompting Large Language Models for Authorship Verification (Hung et al., 2023) offer a new technique named PromptAV. This method utilizes Large Language Models (LLMs) to perform authorship verification effectively and with improved interpretability. Authors claim that the PromptAV, demonstrates improved performance compared to existing state-of-the-art baselines, particularly in scenarios with limited training data. It enhances interpretability by providing intuitive explanations, making it a promising tool for applications in forensic analysis, plagiarism detection, and identifying deceptive content in texts. This approach is meant to address the current limitations of traditional stylometric and deep learning methods, which typically require extensive data and lack explainability (e.g., Bhattacharjee et al., 2023).
@@
The paper T5 meets Tybalt: Author Attribution in Early Modern English Drama Using Large Language Models (Hicke & Mimno, 2023) explores the application of LLMs for authorship identification in Early Modern English drama. The study finds that LLMs, specifically a fine-tuned T5-large model, can accurately predict the author of short passages and outperform traditional baselines like logistic regression, SVM with a linear kernel, and cosine delta. However, the presence of certain authors in the model’s pretraining data introduces biases, leading to occasional confident misattributions of texts. This highlights both the promising potential and the concerning limitations of using LLMs for stylometric analysis in literary studies.
@@
Finally, the paper titled Detecting ChatGPT: A Survey of the State of Detecting ChatGPT-Generated Text (Dhaini et al., 2023) provides an overview of current approaches for identifying text generated by ChatGPT. It highlights the challenges of distinguishing between human-written and machine-generated content, especially given the high fluency and human-like quality of ChatGPT outputs. The survey reviews various datasets specifically created for this detection task, examines different methodologies employed, and discusses qualitative analyses that help identify characteristics unique to ChatGPT-generated text. It also explores the broader implications for domains such as education, law, and science, emphasizing the need for effective MGT detection methods to maintain content integrity.
2.4 LLM detection benchmarks
A non-exhaustive list of datasets designed for MGT detection can be found in Wu et al. (2025). Most of these benchmark datasets are English-only (TuringBench Uchendu et al. (2021), CHEAT Yu et al. (2025), OpenLLMText Chen et al. (2023), GROVER Zellers et al. (2019), TweepFake Fagni et al. (2021), ArguGPT Liu et al. (2023b), MAGE Li et al. (2024a), PAN’s Voight-Kampff Generative AI Detection task Bevendorff et al. (2024)). Others include Spanish (AuTexTification Sarvazyan et al. (2023a)), Chinese (HC3, HC3 Plus Guo et al. (2023), Su et al. (2024)) or rarely they are multilingual (MULTITuDE Macko et al. (2023), M4 Wang et al. (2024b)). The more recent ones are also multi-domain (Macko et al. (2023), Wang et al. (2024b), Bevendorff et al. (2024), Li et al. (2024a), Sarvazyan et al. (2023a), Li et al. (2024a)), and use diverse LLM generators (see especially TuringBench Uchendu et al. (2021), MAGE Li et al. (2024a)), which is particularly challenging to collect for multiple languages. Such multi-generator and multi-domain benchmarks allow one to test generalizability of the classifiers to unseen domains and unseen LLMs, a realistic scenario considering the rate of development of both closed and open-source LLMs. An even more comprehensive overview in terms of domains and languages is given in Macko et al. (2023). It is worth stressing that collecting and generating a well-controlled benchmark with multiple domains, generators, and languages is a considerable endeavour, as is well-known in corpus linguistics Lüdeling & Kytö (2008).
When evaluating MGT detectors, caution is required with regard to training on data external to these benchmarks, as some of the data were collected from other primary sources.
2.5 LLM detection methods
We refer to the Overview of the “Voight-Kampff” Generative AI Authorship Verification Task at PAN and ELOQUENT 2024 by Bevendorff et al. (2024) as a recent benchmark of available methods. The submissions included mainly (i) perplexity-based systems, (ii) term-based systems (iii) and ensembles of both. The first ones rely mainly on the perplexities of a set of known LLMs (which is a limitation on its own) used as features for a classifier. The second class uses fine-tuned classifiers (often neural ones such as modified versions of BERT or their ensembles) with word embeddings or linguistic (stylometric) features; some other involve fine-tuned generative LLMs (regarded as unreliable by Wu et al., 2025). Several proposals, e.g., Guo et al. (2024b), Miralles et al. (2024), Yadagiri et al. (2024), Guo et al. (2024a), made advantage of various sets stylometric features, while others found it useful to augment data or simply expand the training dataset. A noteworthy example is Lorenz et al. (2024), whose SVM classifier based on tf-idf features (sic!) by was ranked third beating all neural baselines and most of the neural-based competitors, mostly thanks to its robustness to the many obfuscation strategies designed by the Task’s authors. We regard our boosted trees classifier to follow a similar simplistic, inexpensive but effective design, often overlooked in recent research.
Among many other methods reviewed by Wu et al. (2025) or Crothers et al. (2023) that were not represented in the PAN’s task are the logits-based statistics. This family encompasses primarily zero-shot methods which, however, require access to a surrogate (often weaker) language model or, ideally, to the source LLM – hence they are dubbed white-box methods – to obtain its raw outputs that in turn allow to determine the likelihood of a text being generated by the LLM. The black-box statistical methods, on the other hand, do not require such access. Instead, given an original text they machine-regenerate it and subsequently compare these versions to obtain a similarity score.
Lastly, a whole strain of research has been MGT watermarking, wherein an imperceptible signature is embedded in the generated texts either by insertion of modified training samples (to defend against unauthorised LLM fine-tuning), manipulating the logits output distribution or token sampling, or character- and word-level replacements at the post-processing stage (especially useful when using black-box models). Sadasivan et al. (2025), Lu et al. (2024) report successful attack strategies on some of such watermarking schemes by iterative paraphrasing or probing a watermarked LLM to infer the signatures. In the present paper, we assume that the adversary does not use a watermarked LLM.
2.6 Robustness of LLM detection
A number of issues can degrade the performance of MGT detection. Wu et al. (2025) divided them into out-of-distribution challenges and attacks.
The first type includes detection across domains (i.e., usually different text types or genres involving different vocabulary, style, topics, and overall distribution of textual features), across languages (but also including texts written by non-native speakers), and across LLMs (i.e., generally detection of LLMs not available during the detector’s training). In the latter case, there are some reports (Sarvazyan et al., 2023b) that supervised MGT detectors tend to generalise well across LLM scales but less so across their model families. For neural detectors, therefore, incorporating MGT from various sources is recommended, since an additional fine-tuning – even on small samples – can effectively alleviate this issue. In our study, no data external to the dataset described in Sec. 3.1.1 is used for training, and we do not tackle the issue of cross-language detection (reported as challenging by Bevendorff et al., 2024). The cross-domain detection is tested on an existing benchmark by Sarvazyan et al. (2023a).
The potential attacks include: paraphrase (where LLM output is subsequently paraphrased by another model in order to change the textual feature distribution of the original MGT; Sadasivan et al., 2025, see, e.g.,), adversarial (involving textual perturbations on the level of characters like various misspelling strategies, Stiff & Johansson, 2022, see , syntax, Bhat & Parthasarathy, 2020, see , or lexis, Crothers et al., 2022, see ), prompt (using complex and varied prompts for MGT, see Guo et al., 2023, Liu et al., 2024b) attacks and models trained specifically to confound existing detectors. These attacks affect MGT detectors differently, depending on whether they are watermarking-based (specifically targeted by paraphrase and adversarial attacks), zero-shot or fine-tuned supervised detectors. The latter can effectively defend against some of these attacks by continually expanding training datasets, e.g., with adversarial examples. Notably, Bevendorff et al. (2024) report that in the joint PAN and ELOQUENT detection-obfuscation task, none of the obfuscation submissions managed to beat in terms of their difficulty simple methods such as Unicode obsfuscations or shortening text length. In this study, we perform a one-step paraphrase attack (repeated paraphrasing is possible, as in Sadasivan et al., 2025), but we do not include any attacks in the training data.
We do not cover the issue of mixed texts (human-edited MGT or LLM-edited human-written texts or texts whose separate parts come from either human or machine).